# Continuized Nesterov Acceleration for Non-Convex Optimization

### v1

@@ 4.1 Improved Factor in Convergence Rate
for some martingale $(M_{t})_{t\geq 0}$ such that $\mathbb{E}\left[M_{t}\right]=0$, $\forall t\in{\mathbb{R}}_{+}$. Evaluating (36) at $t=T_{k}$ and taking expectation, by Theorem 6 we directly deduce the following result.
@@
Comments   As already mentioned, the continuized framework has been applied to strongly quasar convex functions in the specific case of deterministic gradients, i.e. $\rho=1$ in Assumption (SGCρ). The result of Theorem 10 is to compare with (wang2023continuizedaccelerationquasarconvex)
@@ 5 Non-Smooth Lyapunov Functions for Relaxing the Uniqueness Assumption
Because of their definition, strongly quasar convex functions have a unique minimizer. A way to relax this limitation is to use the orthogonal projection operator onto $\mathcal{X}^{\ast}$. It has been successfully used in the case of convex functions (hyppo; aujol2024strong), achieving near-optimal bounds for Nesterov momentum in a setting where former works assumed uniqueness of the minimizer.
@@ 5.1 A Technical Challenge: $\pi(\cdot)$ is not $C^{1}$
$dx_{t}=\zeta(x_{t})dt+\int_{\Xi}G(x_{t^{-}},\xi)dN(t,\xi),$
where $\zeta:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ is a Lipschitz function and $G:{\mathbb{R}}^{d}\times\Xi\to{\mathbb{R}}^{d}$ is measurable. Let $\varphi:{\mathbb{R}}^{d}\to{\mathbb{R}}$ and denote $\phi(t)=\varphi(x_{t})$. If either $\varphi$ is a continuous function such that its directional derivatives exist at all points, or $\phi$ is absolutely continuous, then there exists $\mathbb{A}\subset{\mathbb{R}}_{+}$ of measure zero such that the derivative of $\phi$, denoted $\dot{\phi}$, exists on ${\mathbb{R}}_{+}\backslash\mathbb{A}$. Moreover, if
@@
$\begin{aligned} \sum_{k=0}^{+\infty}\left(\phi(t\wedge T_{k+1})-\phi(t\wedge T_{k+1}^{-})\right)&=\sum_{k=0}^{+\infty}\left(\varphi(x_{{t\wedge T_{k+1}}^{-}}+G(x_{{t\wedge T_{k+1}}^{-}},\xi_{k+1}))-\varphi(x_{{t\wedge T_{k+1}}^{-}})\right)\\ &=\int_{[0,t]}(\varphi(x_{s^{-}}+G(x_{s^{-}},\xi))-\varphi(x_{s^{-}}))dN(s,\xi).\end{aligned}$ (42)
Sum (II)    The sum (II) is composed of $\phi(t\wedge T_{k+1}^{-})-\phi(t\wedge T_{k})$ terms. If $\phi$ was $C^{1}$, we could directly integrate on $\left[t\wedge T_{k},t\wedge T_{k+1}\right)$. Recall that we have either $\varphi$ is a continuous function such that its directional derivatives exist at all points, or $\phi$ is absolutely continuous. We split the proof between these two cases.
Case 1: $\phi$ is absolutely continuous   Absolute continuity of $\phi$ is equivalent to $\phi$ having a derivative $\dot{\phi}$ almost everywhere, which is Lebesgue integrable and verify the fundamental theorem of analysis, namely for $0\leq u\leq t$
$\phi(t)-\phi(u)=\int_{u}^{t}\dot{\phi}(s)ds.$ (43)
Sum (II) thus rewrites
@@
Conclusion for Case 1   Combining (42) and (44), we obtain
$\phi(t)$ $=\phi(0)+\int_{0}^{t}\dot{\phi}(s)ds+\int_{[0,t]}(\varphi(x_{s^{-}}+G(x_{s^{-}},\xi))-\varphi(x_{s^{-}}))dN(s,\xi)$
@@
$+\int_{[0,t]}\int_{\Xi}(\varphi(x_{s^{-}}+G(x_{s^{-}},\xi))-\varphi(x_{s^{-}}))(dN(s,\xi)-dsd\mathcal{P}(\xi))$
@@
Case 2: $\varphi$ is continuous and directionally differentiable everywhere
This case requires a special care. Indeed, without absolute continuity, the current assumption on $\varphi$ are not sufficient to ensure the fundamental theorem of analysis with Lebesgue’s integral, such that we cannot directly write (43). Fortunately, there exists a generalization of the Riemann integral, named Henstock-Kurzweil integral, that allows to obtain a fundamental theorem of analysis of the form (43) under more general assumptions.
Lemma 19
Under the same setting as for Theorem 18, in the case where $\varphi$ is continuous and directionally differentiable everywhere, we have that the derivative of $\phi$, named $\dot{\phi}$, exists except at a countable set $\mathbb{A}$, and
$\begin{aligned} \sum_{k=0}^{+\infty}\left(\phi(t\wedge T_{k+1}^{-})-\phi(t\wedge T_{k})\right)=\sum_{k=0}^{+\infty}\int_{t\wedge T_{i}}^{t\wedge T_{i+1}}\dot{\phi}(s)ds_{HK}=\int_{0}^{t}\dot{\phi}(s)ds_{HK},\end{aligned}$ (45)
where $\int ds_{HK}$ denotes the Henstock-Kurzweil integral.
The proof of Lemma 19 is deferred in Appendix D.1, and see Appendix E.1 for a brief introduction on this integral. Hopefully, this integral shares most of the basic properties of Riemann and Lebesgue integrals, such as linearity or non-negativity, which we use next.
Conclusion for case 2   Combining (42) and (45), we have
$\begin{aligned} \phi(t)&=\phi(0)+\int_{0}^{t}\dot{\phi}(s)ds_{HK}+\int_{[0,t]}(\varphi(x_{s^{-}}+G(x_{s^{-}},\xi))-\varphi(x_{s^{-}}))dN(s,\xi)\\ &=\phi(0)+\int_{0}^{t}\dot{\phi}(s)ds_{HK}+\int_{[0,t]}\mathbb{E}_{\xi}[(\varphi(x_{s}+G(x_{s},\xi))-\varphi(x_{s}))]ds\\ &+\int_{[0,t]}\int_{\Xi}(\varphi(x_{s^{-}}+G(x_{s^{-}},\xi))-\varphi(x_{s^{-}}))(dN(s,\xi)-dsd\mathcal{P}(\xi))\\ &=\phi(0)+\int_{0}^{t}\dot{\phi}(s)ds_{HK}+\int_{[0,t]}\mathbb{E}_{\xi}[(\varphi(x_{s}+G(x_{s},\xi))-\varphi(x_{s}))]ds+M_{t},\end{aligned}$ (46)
where $M_{t}$ is a martingale. Now, as a function that is Lebesgue integrable is Henstock–Kurzweil integrable with same integral value (bartle2001modern, Paragraph 1.10), we have
$\int_{[0,t]}\mathbb{E}_{\xi}[(\varphi(x_{s}+G(x_{s},\xi))-\varphi(x_{s}))]ds=\int_{[0,t]}\mathbb{E}_{\xi}[(\varphi(x_{s}+G(x_{s},\xi))-\varphi(x_{s}))]ds_{HK}.$
By linearity (bartle2001modern, Theorem 3.1), (46) becomes
$\phi(t)=\phi(0)+\int_{0}^{t}\left(\dot{\phi}(s)+\mathbb{E}_{\xi}[(\varphi(x_{s}+G(x_{s},\xi))-\varphi(x_{s}))]\right)ds_{HK}+M_{t}$
This implies that, if we can ensure $\dot{\phi}(s)+\mathbb{E}_{\xi}[(\varphi(x_{s}+G(x_{s},\xi))-\varphi(x_{s}))]\leq 0$, then by non-negativity of the Henstock–Kurzweil integral (bartle2001modern, Theorem 3.2), we have
$\phi(t)\leq\phi(0)+M_{t},$
Proposition 18 gives a tool for Lyapunov analysis in the case of non-$C^{1}$ Lyapunov functions. Intuitively, it indicates that we still have a stochastic derivation rule to ensure $\mathbb{E}\left[\phi(t)\right]\leq\phi(0)$. Having to show $\dot{\phi}(s)+\mathbb{E}_{\xi}[(\varphi(x_{s}+G(x_{s},\xi))-\varphi(x_{s}))]\leq 0$ corresponds to having to show $\left<\nabla\varphi(x_{s}),\zeta(x_{s})\right>+\mathbb{E}_{\xi}\varphi(x_{s}+G(x_{s}))-\varphi(x_{s})\leq 0$ in the differentiable case, see e.g. the proof of Proposition 7.
@@
Absolute continuity can be difficult to verify. More conveniently, functions $g:{\mathbb{R}}\to{\mathbb{R}}^{d}$ that are $C^{1}$ or Lipschitz are automatically absolutely continuous. In many applications, however, Lyapunov functions typically involve the function we want to minimize, which may fail to be $C^{1}$ or Lipschitz—for instance in non-smooth composite optimization. The alternative condition of Proposition 18, that is that $\varphi$ is a continuous function such that its directional derivatives exist at every points, may thus provide a promising avenue.
@@ 5.2 Convergence Result
This Lyapunov function is not $C^{1}$, because of the projection. However it is continuous and directionally differentiable, as the partial derivatives of $\pi(\cdot)$ exist (bonnans1998sensitivity, Theorem 7.2). We could also deduce from the convexity of $\mathcal{X}^{\ast}$ that $\pi(\cdot)$ is $1$-Lipschitz, which would allow to leverage absolute continuity. Both viewpoints would allow us to apply Proposition 18, but we chose the first one since it yields a slightly simpler proof.
@@ 6.1 Proof of Theorem 10
Proof Let $\overline{x}_{t}=(t,x_{t},z_{t})$. It satisfies $d\overline{x}_{t}=\zeta(\overline{x}_{t})dt+G(\overline{x}_{t},\xi)dN(t,\xi)$, where
@@
Using $\tau,\mu$-strong quasar convexity, we get
@@ 6.2 Proof of Theorem 21
Let $\overline{x}_{t}=(t,x_{t},z_{t})$. It satisfies $d\overline{x}_{t}=\zeta(\overline{x}_{t})dt+G(\overline{x}_{t},\xi)dN(t,\xi)$, where
@@
The property of the orthogonal projection onto a convex set has been studied extensively (ZARANTONELLO1971237). $\mathcal{X}^{\ast}$ being a closed convex set with $C^{2}$ boundary ensures $\pi$ is well defined on ${\mathbb{R}}^{d}$ (holmes1973smoothness; fitzpatrick1982differentiability), such that its partial derivatives exist (bonnans1998sensitivity, Theorem 7.2). Noting $\phi(t):=\varphi(\overline{x}_{t})$, we apply Proposition 18, such that $\dot{\phi}$ the derivative of $\phi$ exists on ${\mathbb{R}}_{+}\backslash\mathbb{A}$, where $\mathbb{A}$ is a set of null measure. In particular on ${\mathbb{R}}_{+}\backslash\mathbb{A}$, $(x_{t},z_{t})_{t\geq 0}$ are $C^{1}$ and verify
@@
Acknowledgments and Disclosure of Funding
This work was supported by the ANR project PEPR PDE-AI.
@@ D.1 Proof of Lemma 19
For $s\in(t\wedge T_{i},t\wedge T_{i+1})$, we have
$\dot{x}_{t}=\zeta(x_{t}),$ (129)
such that $x_{s}$ is $C^{1}$ on $(t\wedge T_{i},t\wedge T_{i+1})$. Noting $\mathbb{T}:=\{T_{n}\}_{n\in{\mathbb{N}}}$, we have that $(x_{t})_{t\geq 0}$ is $C^{1}$ verifying (129) on ${\mathbb{R}}_{+}\backslash\mathbb{T}$. Because $\varphi$ is directionally differentiable, we then have $\phi:t\mapsto\varphi(x_{t})$ is left and right differentiable on this interval. Formally, for all negative sequences $\{h_{i}\}_{i\in{\mathbb{N}}}$ and a positive sequence $\{h^{\prime}_{i}\}_{i\in{\mathbb{N}}}$, both with zero as limit, we have that
$\lim_{i\to+\infty}\frac{\phi(x+h_{i})-\phi(x)}{h_{i}}=\phi^{-}(x),\quad\lim_{i\to+\infty}\frac{\phi(x+h^{\prime}_{i})-\phi(x)}{h^{\prime}_{i}}=\phi^{+}(x).$ (130)
We then use the following result, already used by (hyppo) in order to relax an assumption of unique minimizer in a convex setting.
Corollary 28 ((young1914note, Paragraph 6))
[…]
$\lim_{i\to+\infty}\frac{g(x+h_{i})-g(x)}{h_{i}}=\lim_{i\to+\infty}\frac{g(x+h^{\prime}_{i})-g(x)}{h^{\prime}_{i}},$
i.e. the exists a symmetric derivative.
As $\phi$ restricted to $t\wedge T_{i},t\wedge T_{i+1})$ is continuous, we can apply the above result. Because of (130), it implies $\phi^{-}=\phi^{+}$ except at a countable set, that we denote $\mathbf{S}_{i}$. We then have that $\phi$ is derivable on $(t\wedge T_{i},t\wedge T_{i+1})\penalty 10000\ \backslash\penalty 10000\ \mathbf{S_{i}}$, and we note $\dot{\phi}$ its derivative. The fact that $\phi$ is continuous on $t\wedge T_{i},t\wedge T_{i+1})$ , whose derivative exists except on a countable set, makes it Henstock–Kurzweil integrable on this interval. In particular, one has
$\phi(t\wedge T_{i+1}^{-})=\phi(t\wedge T_{i})+\int_{t\wedge T_{i}}^{t\wedge T_{i+1}^{-}}\dot{\phi}(s)ds_{HK}.$ (131)
We replicate the argument on each interval $t\wedge T_{j},t\wedge T_{j+1})$, $j\in{\mathbb{N}}$, such that the derivative of $\phi$ exists on those interval except at a countable set $\mathbf{S_{j}}$. Summing over all the interval, we obtain that $\dot{\phi}$ exists on $\mathbb{R}$$_{+}$ except at $\mathbb{A}:=\bigcup_{j\in{\mathbb{N}}}\mathbf{S_{j}}\bigcup\mathbb{T}$, which is a countable set. Summing over $k\in{\mathbb{N}}$, we obtain
$\sum_{k=0}^{+\infty}\left(\phi(t\wedge T_{k+1}^{-})-\phi(t\wedge T_{k})\right)=\sum_{k=0}^{+\infty}\int_{t\wedge T_{i}}^{t\wedge T_{i+1}}\dot{\phi}(s)ds_{HK}=\int_{0}^{t}\dot{\phi}(s)ds_{HK}.$
D.2 Proof of Lemma 1
@@ Appendix E Background
In this section, we collect some known properties about Henstock–Kurzweil integrals (Section E.1), and justify the link between two formulations of ODEs associated to momentum algorithm (Section E.3).
E.1 Some Properties of the Henstock–Kurzweil Integral
In this section, as we use it in our Section 5, we give a brief introduction and we state, without proof, some properties of the Henstock–Kurzweil integral, sometimes referred to as the generalized integral. The theory of this integral has been independently developed by Kurzweil (kurzweil1957generalized) and Henstock (Henstock1963). For a detailed presentation, we refer to (bartle2001modern).
In the whole section, $I=[a,b]$ denotes an interval of $\mathbb{R}$}.
Definition 29
$(I_{k},\nu_{k})_{k=1}^{n}$ is a tagged partition of $I$ if
  1. 1.
    $(I_{k})_{k=1}^{n}$ is a partition of intervals of $I$.
  2. 2.
    $\nu_{k}\in I_{k}$, $\forall k\in\{1,\dots,n\}$.
A tagged partition gives a partition of $I$ and, and for each interval of the partition, a point in this interval. This object is also involved in the classical Riemann integral. The difference here will be the use of gauge functions.
Definition 30
An application $\delta:I\to(0,+\infty)$ is a gauge function.
Intuitively, we will use gauge functions to control the lengths of the intervals in the partitions of $I$. This is formally done with the idea of $\delta$-fine partition.
Definition 31
For a gauge $\delta:I\to(0,+\infty)$, a tagged partition $(I_{k},\nu_{k})_{k=1}^{n}$ is $\delta$-fine if
$I_{k}\subset[\nu_{k}-\delta(\nu_{k}),\nu_{k}+\delta(\nu_{k})].$
For a defined gauge function, such a partition always exists (cousin1895fonctions). Now, we can define what it means to be Kurzweil-integrable.
Definition 32
A function $f:I\to{\mathbb{R}}$ is Kurzweil-Henstock integrable with real value denoted $\int_{I}f(s)ds_{HK}$ if for all $\epsilon>0$, there exists a gauge $\delta_{\varepsilon}$ such that for all tagged partition $(I_{k},\nu_{k})_{k=1}^{n}$ that is $\delta_{\varepsilon}$-fine, the Riemann-Kurzweil-Henstock sum verifies
$\left|\sum_{i=k}^{n}|I_{k}|f(\nu_{k})-\int_{I}f(x)ds_{HK}\right|\leq\varepsilon,$
with $|I_{k}|$ the length of $I_{k}$.
A function that is Riemann or Lebesgue integrable is also Kurzweil-Henstock integrable, and in this case their value coincide. Kurzweil-Henstock integrals verify similar useful properties as for the classical integrals.
Proposition 33
Let $f,g:I\to{\mathbb{R}}$ be Kurzweil-Henstock integrable. Then
  1. 1.
    (Linearity, (bartle2001modern, Theorem 3.1)) $\forall a,b\in{\mathbb{R}}$, $\int_{I}(af+bg)d_{HK}=a\int_{I}fd_{HK}+b\int_{I}gd_{HK}$.
  2. 2.
    (non-negativity, (bartle2001modern, Theorem 3.2)), $f(x)\geq 0$, $\forall x\in I\Rightarrow\int_{I}fd_{HK}\geq 0$.
  3. 3.
    (Additivity, (bartle2001modern, Theorem 3.7)) For $a\leq c\leq b$, $\int_{a}^{c}fd_{HK}+\int_{c}^{b}fd_{HK}=\int_{a}^{c}fd_{HK}$.
  4. 4.
    (Hake’s Theorem, (bartle2001modern, Theorem 12.8)) $\lim_{c\to b^{-}}\int_{a}^{c}fd_{HK}=\int_{a}^{b}fd_{HK}$.
We end this section by stating the remarkable property of this type of integral: the fundamental theorem of calculus holds under more relaxed assumptions than compared to Riemann and Lebesgue’s integrals.
Theorem 34
Let $F:[a,b]\to R$. If either
  1. 1.
    $F^{\prime}(x)=f(x)$ for all $x\in I$.
  2. 2.
    If $F$ is continuous and $F^{\prime}(x)=f(x)$ for all $x\in I$ except on a countable set.
then $f$ is Kurzweil-Henstock integrable and
$F(b)=F(a)+\int_{I}f(x)dx_{HK}.$
This result are Theorem 4.5 and 4.7 in (bartle2001modern). The second point is the one we use in our proof of Proposition 18. Importantly, these assumptions are too weak to ensure a similar result when using the Lebesgue integral. In particular, one can construct examples of continuous functions, derivable almost everywhere such that the fundamental theorem of calculus does not hold (bartle2001modern, Theorem 4.17).
E.2 The different forms of Momentum

### v2

@@ 4.1 Improved Factor in Convergence Rate
for some martingale $(M_{t})_{t\geq 0}$ such that $\mathbb{E}\left[M_{t}\right]=0$, $\forall t\in{\mathbb{R}}_{+}$. Evaluating (36) at $t=T_{k}$ and taking expectation, by Theorem 6 and Proposition 5 we directly deduce the following result.
@@
Comments   As already mentioned, the continuized framework has been applied to strongly quasar-convex functions in the specific case of deterministic gradients, i.e. $\rho=1$ in Assumption (SGCρ). The result of Theorem 10 is to compare with (wang2023continuizedaccelerationquasarconvex)
@@ 5 Non-Smooth Lyapunov Functions for Relaxing the Uniqueness Assumption
Because of their definition, strongly quasar-convex functions have a unique minimizer. A way to relax this limitation is to use the orthogonal projection operator onto $\mathcal{X}^{\ast}$. It has been successfully used in the case of convex functions (hyppo; aujol2024strong), achieving near-optimal bounds for Nesterov momentum in a setting where former works assumed uniqueness of the minimizer.
@@ 5.1 A Technical Challenge: $\pi(\cdot)$ is not $C^{1}$
$dx_{t}=\zeta(x_{t})dt+\int_{\Xi}G(x_{t^{-}},\xi)N(dt,d\xi),$
where $\zeta:{\mathbb{R}}^{d}\to{\mathbb{R}}^{d}$ is a locally Lipschitz function and $G:{\mathbb{R}}^{d}\times\Xi\to{\mathbb{R}}^{d}$ is measurable. Let $\varphi:{\mathbb{R}}^{d}\to{\mathbb{R}}$ be locally Lipschitz. Then, denoting $\phi(t)=\varphi(x_{t})$, there exists $\mathbb{A}\subset{\mathbb{R}}_{+}$ of zero measure such that the derivative of $\phi$, denoted by $\dot{\phi}$, exists on ${\mathbb{R}}_{+}\backslash\mathbb{A}$. Moreover, if
@@
$\begin{aligned} \sum_{k=0}^{\infty}\phi(t\wedge T_{k+1})-\phi(t\wedge T_{k+1}^{-})&=\sum_{k\geq 0:\,T_{k+1}\leq t}\phi(T_{k+1})-\phi(T_{k+1}^{-})\\ &=\sum_{k\geq 0:\,T_{k+1}\leq t}\varphi(x_{T_{k+1}^{-}}+G(x_{T_{k+1}^{-}},\xi_{k+1}))-\varphi(x_{T_{k+1}^{-}})\\ &=\int_{0}^{t}\int_{\Xi}\varphi(x_{s-}+G(x_{s-},\xi))-\varphi(x_{s-})\,N(ds,d\xi).\end{aligned}$ (42)
Sum (II)    The sum (II) is composed of $\phi(t\wedge T_{k+1}^{-})-\phi(t\wedge T_{k})$ terms. For $s\in(t\wedge T_{k},t\wedge T_{k+1})$, we have
$\dot{x}_{t}=\zeta(x_{t}),$ (43)
such that $x_{s}$ is $C^{1}$ on $(t\wedge T_{k},t\wedge T_{k+1})$. Together with the fact that $\varphi$ is locally Lipschitz, it induces that $\phi(t)=\varphi(x_{t})$ is locally Lipschitz on $(t\wedge T_{k},t\wedge T_{k+1})$. Then, $\phi$ is locally absolutely continuous on $(t\wedge T_{k},t\wedge T_{k+1})$. This means that its derivative $\dot{\phi}$ exists almost everywhere on $(t\wedge T_{k},t\wedge T_{k+1})$, and that we have
$\phi(t\wedge T_{k+1}^{-})-\phi(t\wedge T_{k})=\int_{t\wedge T_{k}}^{t\wedge T_{k+1}}\dot{\phi}(s)ds,$
where we used that $\phi$ is right-continuous at $T_{k}$ because $(x_{t})_{t\geq 0}$ is càdlàg. Applying the same reasoning on each intervals $(t\wedge T_{k},t\wedge T_{k+1})$, $k\in{\mathbb{N}}$, we have that the derivative of $\phi$ exists almost everywhere, and that Sum (II) rewrites
@@
Conclusion   Combining (42) and (44), we obtain
$\phi(t)$ $=\phi(0)+\int_{0}^{t}\dot{\phi}(s)ds+\int_{[0,t]}(\varphi(x_{s^{-}}+G(x_{s^{-}},\xi))-\varphi(x_{s^{-}}))N(ds,d\xi)$
@@
$+\int_{[0,t]}\int_{\Xi}(\varphi(x_{s^{-}}+G(x_{s^{-}},\xi))-\varphi(x_{s^{-}}))(N(ds,d\xi)-ds\mathcal{P}(d\xi))$
@@
Compared to Proposition 3, Proposition 18 provides a Lyapunov analysis framework in the case of non-$C^{1}$ Lyapunov functions. Precisely, it relaxes the $C^{1}$ assumption to allow locally Lipschitz functions. Intuitively, it indicates that even with non-differentiable Lyapunov function, we still have a rule to ensure $\mathbb{E}\left[\phi(t)\right]\leq\phi(0)$. In this case, having to show $\dot{\phi}(s)+\mathbb{E}_{\xi}[(\varphi(x_{s}+G(x_{s},\xi))-\varphi(x_{s}))]\leq 0$ plays the same role as establishing $\left<\nabla\varphi(x_{s}),\zeta(x_{s})\right>+\mathbb{E}_{\xi}\varphi(x_{s}+G(x_{s}))-\varphi(x_{s})\leq 0$ in the differentiable case, see e.g. the proof of Proposition 7.
@@
As the function $f$ we minimize is typically involved in Lyapunov functions, Proposition 18 imposes $f$ to be locally Lipschitz. This is a natural assumption in non-smooth optimization, as it ensures that the slopes are locally bounded. However, it implies $f$ is continuous. It might be necessary to carry the continuized analysis, as we want to write $\phi$ as an integral in the proof of Proposition 18.
@@ 5.2 Convergence Result
This Lyapunov function is not $C^{1}$ because of the projection, and we cannot apply Proposition 3. However $\pi(\cdot)$ is Lipschitz, so in particular locally Lipschitz, such that Proposition 18 applies.
@@ 6.1 Proof of Theorem 10
Proof Let $\overline{x}_{t}=(t,x_{t},z_{t})$. It satisfies $d\overline{x}_{t}=\zeta(\overline{x}_{t})dt+G(\overline{x}_{t},\xi)N(dt,d\xi)$, where
@@
Using $\tau,\mu$-strong quasar-convexity, we get
@@ 6.2 Proof of Theorem 20
Let $\overline{x}_{t}=(t,x_{t},z_{t})$. It satisfies $d\overline{x}_{t}=\zeta(\overline{x}_{t})dt+G(\overline{x}_{t},\xi)N(dt,d\xi)$, where
@@
The property of the orthogonal projection onto a convex set has been studied extensively (ZARANTONELLO1971237). $\mathcal{X}^{\ast}$ being a closed convex set with $C^{2}$ boundary ensures $\pi$ is well defined on ${\mathbb{R}}^{d}$ (holmes1973smoothness; fitzpatrick1982differentiability), and in particular, it is $1$-Lipschitz. As $f$ is $C^{1}$, we have that $\varphi$ is locally Lipschitz. Noting $\phi(t):=\varphi(\overline{x}_{t})$, we apply Proposition 18, such that $\dot{\phi}$ the derivative of $\phi$ exists on ${\mathbb{R}}_{+}\backslash\mathbb{A}$, where $\mathbb{A}$ is a set of null measure. In particular on ${\mathbb{R}}_{+}\backslash\mathbb{A}$, $(x_{t},z_{t})_{t\geq 0}$ are $C^{1}$ and verify
@@ Appendix E Background
In this section, we give additional details about different formulations of momentum ODEs and algorithms.
E.1 The Different Forms of Momentum